OpenStudy (anonymous):

solve. 9 over z+4=8 over z-4

OpenStudy (saifoo.khan):

let's see who wins..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[ \frac{9}{z+4} = \frac{8}{z-4} \] \[9z - 36 = 8z + 32\] \[z = 32 +36\] \[x =78\]

OpenStudy (saifoo.khan):

lil answered, but incorrect.

OpenStudy (saifoo.khan):

winner = ishaan.

OpenStudy (saifoo.khan):

wait, that's 68. LOL

OpenStudy (saifoo.khan):

stop eating fries ishaan.

OpenStudy (anonymous):

yay! \(\LaTeX\) Power oh yeah 68 sorry not 78 but 68

OpenStudy (anonymous):

Lol I love fries

OpenStudy (anonymous):

thanks guys. you all are life savers

OpenStudy (saifoo.khan):

uw.

OpenStudy (anonymous):

no problem :)